OpenStudy (anonymous):

If M (3,-5) is the midpoint of A (-3,1) and B(X,Y) find the coordinates of B. Please help me solve this! :)

OpenStudy (angela210793):

XM=(XA+XB)/2--->6=-3+x--->xB=9 YM=(YA+YB)/2--->2=1+yA--->yA=1 B(9;1)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (dumbcow):

The midpoint of 2 points is just the average of the coordinates

OpenStudy (angela210793):

w8 i've done smth wrong

OpenStudy (anonymous):

oh ok

OpenStudy (dumbcow):

i get (9,-11)

OpenStudy (angela210793):

YM=(YA+YB)/2---> -10=1+yB--->yB=-11

OpenStudy (anonymous):

Since M is the midpoint of A, you know that your line should extend in the fourth quadrant. \[\left(\begin{matrix}x2+x1 \\ 2\end{matrix}\right), \left(\begin{matrix}y2+y1 \\ 2\end{matrix}\right)\] is your midpoint formula.

OpenStudy (angela210793):

yea the answer is (9;-11)

OpenStudy (anonymous):

Thank you all very much for the quick and easy help! :)

OpenStudy (dumbcow):

your welcome